# Learning to Compress Contexts for Efficient Knowledge-based Visual Question Answering

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated great zero-shot performance on visual question answering (VQA). However, when it comes to knowledge-based VQA (KB-VQA), MLLMs may lack human commonsense or specialized domain knowledge to answer such questions and require obtaining necessary information from external knowledge sources. Previous works like Retrival-Augmented VQA-v2 (RAVQA-v2) focus on utilizing as much input information, such as image-based textual descriptions and retrieved knowledge, as possible to improve performance, but they all overlook the issue that with the number of input tokens increasing, inference efficiency significantly decreases, which contradicts the demands of practical applications. To address this issue, we propose Retrieval-Augmented MLLM with Compressed Contexts (RACC). RACC learns to compress and aggregate retrieved contexts, from which it generates a compact modulation in the form of Key-Value (KV) cache. This modulation is then used to adapt the downstream frozen MLLM, thereby achieving effective and efficient inference. RACC achieves a state-of-the-art (SOTA) performance of 62.9% on OK-VQA. Moreover, it significantly reduces inference latency by 22.0%-59.7% compared to the prominent RAVQA-v2. Abundant experiments show RACC’s broad applicability. It is compatible with various off-the-shelf MLLMs and can also handle different knowledge sources including textual and multimodal documents.
Introduction
@@
In previous studies of KB-VQA, a line of works (Hu et al. 2023a; Khademi et al. 2023; An et al. 2024) obtains knowledge from very large MLLMs (GPT-4) or LLMs (ChatGPT, GPT-3) that have been pretrained on extensive corpora. However, the static knowledge in these models can be out-of-date and potentially lead to unreliable knowledge due to hallucinations, particularly in specific domains. Another line of research retrieves knowledge from external knowledge sources, such as knowledge graphs (Speer, Chin, and Havasi 2017), documents (Luo et al. 2021) etc., which are cheaper, more reliable and up-to-date, better aligning with the needs of real-world applications. Retrieval Augmented VQA-v2 (RAVQA-v2) represents a significant advancement in this kind of approach, where it concatenates $K$ retrieved documents with each image-question pair and puts them into MLLMs to generate $K$ candidate answers. RAVQA-v2 tackles the KB-VQA problem by performing straightforward retrieval-augmented generation (RAG) on MLLMs. However, it has a notable drawback, i.e. low efficiency during inference. In the inference process of RAVQA-v2, $K$ candidate answers are initially generated based on $K$ retrieved documents, and the final answer is then selected from the $K$ candidates by their joint probabilities, which is undoubtedly very time-consuming and resource-intensive. Moreover, the retrieved documents can be quite long and often contain a lot of redundant information, which can further exacerbate the problem of low inference efficiency.
@@
However, inference efficiency is a key concern in practical applications of MLLMs. Previous KB-VQA works have consistently focused on how to use as much knowledge as possible to improve performance, but they have overlooked the issue that inference efficiency significantly declines as the number of input tokens increases.
@@
Therefore, in this work, we aim to propose a new RAG framework based on MLLMs, which can utilize the information of retrieved contexts in an effective and efficient manner to improve MLLMs’ inference efficiency for KB-VQA.
@@
Furthermore, RAVQA-v2 and many previous works (Luo et al. 2021; Lin and Byrne 2022; Lin et al. 2022) on KB-VQA primarily focus on using textual documents as external knowledge sources, and there has been relatively less research on using multimodal documents as external knowledge sources. However, multimodal documents are a common and important knowledge resource in real-world applications, and Hu et al. (2023b) demonstrates that leveraging information from multimodal documents can provide knowledge for models like T5 (Raffel et al. 2020) in handling KB-VQA tasks. More importantly, MLLMs inherently have the ability to directly comprehend multimodal knowledge, and we believe that exploring the effects of multimodal documents in RAG applications based on MLLMs is of significant importance. However, this area remains unexplored. We hope that our proposed framework can effectively utilize the inherent capabilities of MLLMs to employ both textual and multimodal documents as knowledge sources in a unified, effective, and efficient manner, thereby enhancing the KB-VQA performance of MLLMs.
Therefore, in this paper, we propose RACC, i.e. Retrieval-Augmented MLLMs with Compressed Contexts, an effective and efficient RAG framework for KB-VQA based on MLLMs. Our proposed framework first leverages a frozen hyperMLLM to learn to compress retrieved documents into short soft prompts. Then, we design an elaborate aggregator module to aggregate compressed prompts. Finally, a set of Multi-Layer Perceptrons (MLPs) is used to generate a compact modulation in the form of Key-Value (KV) cache to adapt the downstream frozen baseMLLM. With the compact modulation, the baseMLLM can utilize the information in the retrieved documents in a highly efficient manner. Our contributions can be summarized as follows:
@@
  • RACC achieves excellent performance comparable to many competitive baselines on two KB-VQA datasets at a very low cost, reaching a state-of-the-art (SOTA) performance of 62.9% on the OK-VQA dataset.
@@
  • We are the first to explore how to perform retrieval-augmented generation on MLLMs in a highly efficient manner. For time efficiency, RACC can save 22.0-59.7% of inference latency compared to RAVQA-v2. For space efficiency, RACC supports pre-saving documents that occupy a large storage footprint in the form of compressed prompts to save disk space.
@@
  • Furthermore, abundant experiments demonstrate that RACC can be applied to various off-the-shelf MLLMs, but also can handle different knowledge sources such as textual documents and multimodal documents.
Figure 1: The structural framework of RACC.
Related Work
Multimodal Large Language Models
@@
Multimodal Large Language Models (MLLMs) bridge the gap between text modality and other modalities and unifies the understanding of different modalities. Although recent MLLMs (Alayrac et al. 2022; Li et al. 2023; Dai et al. 2023; Bai et al. 2023; Liu et al. 2023; Hu et al. 2024) have demonstrated excellent zero-shot performance on VQA tasks, they need the assistance of external knowledge sources for questions that require specialized domain knowledge to answer. In this work, we investigate MLLMs of two mainstream architectures: the encoder-decoder MLLMs including BLIP2-FlanT5XL (Li et al. 2023) and InstructBLIP-FlanT5XL (Dai et al. 2023), and the decoder-only models InstructBLIP-Vicuna7B and miniCPM-v2 (Hu et al. 2024).
Knowledge-based Visual Question Answering
KB-VQA is a task that has received a lot of attention in the current era, aiming to use external knowledge to answer image-based questions.
@@
Early KB-VQA works (Luo et al. 2021; Marino et al. 2021) train specialized models with designated knowledge sources such as knowledge graphs (ConceptNet-Speer, Chin, and Havasi 2017), textual document sources (Google Search-Luo et al. 2021; WikiData-Vrandečić and Krötzsch 2014), images sources (Google Image Search), etc.. However, these methods often demonstrate limited performance.
@@
With the emergence of pretrained large language models (LLMs), they have become a focus of research in this field. A series of works utilize very large LLMs like GPT-3 to generate auxiliary knowledge or directly answer the questions (Gui et al. 2021; Lin et al. 2022; Yang et al. 2022; Shao et al. 2023; Hu et al. 2023a), achieving performance breakthroughs. Due to the high costs associated with using very large LLMs like GPT-3, and the fact that the knowledge within them can be outdated and incorrect, another line of work focuses on how to conduct retrieval-augmented generation (RAG) on smaller LLMs with various external knowledge bases for KB-VQA (Gui et al. 2021; Gao et al. 2022; Lin and Byrne 2022; Lin et al. 2024a). REVEAL (Hu et al. 2023b) designs a complex pipeline that utilizes a ViT encoder (Dosovitskiy et al. 2020) and a T5 encoder (Raffel et al. 2020) to extract information from multimodal knowledge sources and generates answers by a T5 decoder, which requires extensive pre-training to learn how to utilize multimodal knowledge. However, methods based on LLMs all face the same problem: they often require converting provided images into textual descriptions such as captions and object tags so that LLMs can understand (Gui et al. 2021; Gao et al. 2022; Shao et al. 2023; Lin et al. 2024a), which may result in loss of critical visual information in the images but also significantly increases the number of input tokens, leading to a notable rise in inference latency.
@@
With the advent of multimodal large language models (MLLMs), the aforementioned problems have been perfectly resolved. Recent works have made new progress by utilizing MLLMs. A line of works proposes to combine MLLMs and LLMs together (Khademi et al. 2023; Xenos et al. 2023; An et al. 2024). MM-Reasoner (Khademi et al. 2023) leverages vision APIs and rationales generated by GPT-4 to fine-tune MLLMs such as Flamingo. RAVQA-v2 (Lin et al. 2024a) builds a simple RAG framework on top of pure MLLMs. However, as mentioned earlier, it suffers from low efficiency during the inference stage.
Prompt Compression
@@
Given the inherent redundancy in natural language, prompt compression methods have been extensively studied to improve the efficiency of LLM inference. Prompt compression can be categorized into task-aware and task-agnostic methods. Since the generation of compressed prompts that perform well across diverse tasks is particularly challenging, we focus on the task-aware prompt compression paradigm. An important line of work Jiang et al. (2023a); Pan et al. (2024); Jiang et al. (2023b) estimates the importance of the tokens within the original prompts by the information-based metric etc. and removes redundant tokens. Xu, Shi, and Choi (2024) trains a compressor model, which cuts redundant tokens in the passage based on the question.
@@
In addition to the above methods, which detect and remove inherent redundant tokens in long contexts at the natural language level, a series of works aims at compressing long contexts into parameters, leveraging the capability of LLMs to implicitly eliminate redundant information within long contexts. Mu, Li, and Goodman (2024) supposes that each prompt is composed of a task instruction part and a content part, and finetunes LLMs to compress the task instruction part into several gist tokens. Chevalier et al. (2023); Wang, Ma, and Cai (2024); Tack et al. (2024) learns to compress long contexts into compact summary vectors, parameters of a Lora-module and KV Cache, respectively.
@@
RACC-hetero BLIP2-Vicuna7B GS 60.67
RACC-hetero InstructBLIP-Vicuna7B GS 62.91
@@
Table 1: Model Performance on the OK-VQA dataset. Knowledge source abbreviations: C: ConceptNet; CC: CC12M; V2: VQA-2; W: Wikipedia; WD: WikiData; WIT: Wikipedia Image-Text; GS: Google Search; GI: Google Images; ICE: In-context Examples; PNP: Plug-and-Play VQA captioner (Tiong et al. 2022). In RACC-homo, the hyperMLLM and baseMLLM share the same structures and weights, while in RACC-hetero, they differ in either structure or weights. In the last two rows of results of the RACC-hetero, the hyperMLLM used is InstructBLIP-FlanT5XL.
Methods
Problem Setup
A typical VQA dataset can be divided into three components: images, questions, and answers, which can be represented using the notation $\{v,q,a\}_{n}$.
@@ Methods
Following (Lin et al. 2024a), we consider a realistic scenario of KB-VQA: an MLLM that takes an image $v_{i}$ and its related question $q_{i}$ as input, where the knowledge required to answer the question is supplied by external knowledge sources. In this paper, we study two common and essential knowledge sources in real-world applications: documents, including multimodal documents and textual documents.
@@
We utilize an off-the-shelf frozen multimodal retriever to retrieve $K$ documents from the given knowledge source conditioned on the provided image and question. The $K$ retrieved documents is denoted as $\{d_{i}\}^{K}=\{d^{1}_{i},d^{2}_{i},\ldots,d^{K}_{i}\}$. The output confidence scores corresponding to the K documents are denoted as $\{p_{i}\}^{K}=\{p^{1}_{i},p^{2}_{i},\ldots,p^{K}_{i}\}$. The MLLM needs to leverage these relevant documents to provide the correct answer to the question $q_{i}$ based on $v_{i}$.
Our Proposed Framework: RACC
Building on the above setup, we propose RACC, i.e. Retrieval-Augmented MLLM with Compress Contexts. In this section, we first delineate our framework’s overall workflow and then discuss the strategies we have designed to optimize it in accordance with the task-specific characteristics. The framework structure of RACC is depicted in Figure 1.
Learning to compress contexts by the hyperMLLM.
@@
The first step in our framework is to compress the retrieved document into soft prompts of a specified length. Phang et al. (2023); Tack et al. (2024) introduces the idea of amortized-based meta-learning into the online learning task, which brings us inspiration. We utilize MLLMs based on the encoder-decoder architecture (such as BLIP2-FlanT5XL (Li et al. 2023), InstructBLIP-FlanT5XL (Dai et al. 2023), etc.) with a set of learnable prompts for the decoder to compress the input information. We refer to the frozen MLLM used to compress retrieved documents as hyperMLLM and denote it as $g_{hyper}$. The compressing process is denoted as follows:
$\vspace{-1em}\theta^{k}_{i}=g_{hyper}(d^{k}_{i},\theta_{d}),$ (1)
where $\theta^{k}_{i}$ denotes the compressed prompts of the document $d^{k}_{i}$ and $\theta_{d}$ represents the predefined learnable prompts for compressing retrieved documents.
We also compress the image-question pairs in a similar manner. The key differences are that we first process the image and question independently using the hyperMLLM, and subsequently, we concatenate them for an additional round of compression. The reason behind this design choice will be elaborated in the following section. Denoting the predefined learnable prompts for compressing the image-question pairs as $\theta_{vq}$, the compression process is as follows:
$\begin{split}&\theta_{v_{i}}=g_{hyper}(v_{i},\theta_{vq}),\\ &\theta_{q_{i}}=g_{hyper}(q_{i},\theta_{vq}),\\ &\theta_{vq_{i}}=g_{hyper}(\textsc{concat}(v_{i},q_{i}),\theta_{vq})\end{split}$ (2)

### v2

@@ Abstract
Multimodal large language models (MLLMs) have demonstrated great performance on visual question answering (VQA). When it comes to knowledge-based Visual Question Answering (KB-VQA), MLLMs may lack the specialized domain knowledge needed to answer questions, necessitating the retrieval of necessary information from external knowledge sources. Previous works like Retrival-Augmented VQA-v2 (RAVQA-v2) focus on utilizing as much input information, such as image-based textual descriptions and retrieved knowledge, as possible to improve performance, but they all overlook the issue that with the number of input tokens increasing, inference efficiency significantly decreases, which contradicts the demands of practical applications. To address this issue, we propose Retrieval-Augmented MLLMs with Compressed Contexts (RACC). RACC learns to compress and aggregate retrieved knowledge for a given image-question pair, generating a compact modulation in the form of Key-Value (KV) cache to adapt the downstream frozen MLLM, thereby achieving effective and efficient inference. RACC achieves a state-of-the-art (SOTA) performance of 63.92% on OK-VQA. Moreover, it significantly reduces inference latency by 22.0%-59.7% compared to the prominent RAVQA-v2. Abundant experiments show RACC’s broad applicability. It is compatible with various off-the-shelf MLLMs and can also handle different knowledge sources including textual and multimodal documents.
Machine Learning, ICML
1 Introduction
@@
In previous studies of KB-VQA, a line of works (Hu et al., 2023a; Khademi et al., 2023; An et al., 2024) utilize knowledge from very large MLLMs (GPT-4) or LLMs (ChatGPT, GPT-3) to aid in answering question. However, the static knowledge in such models may become out-of-date, and the models may generate incorrect content due to hallucinations, particularly in specific domains. What’s more, the high cost associated with using these models is also a notable downside. Another line of research which retrieves knowledge from updateable external knowledge sources, such as knowledge graphs (Speer et al., 2017), documents (Luo et al., 2021), etc., is more reliable and better aligns with the needs of real-world applications.
@@
Retrieval Augmented VQA-v2 (RAVQA-v2) tackles the problem of KB-VQA by performing straightforward retrieval-augmented generation (RAG) on MLLMs. However, it has a notable shortcoming, i.e. low efficiency during inference. In the inference process of RAVQA-v2, the $K$ retrieved documents are first concatenated with the image-question pair and inputted into the MLLM to obtain $K$ candidate answers, then the final answer is selected from the $K$ candidates by their joint probabilities. The process is undoubtedly very time-consuming and resource-intensive. Moreover, the retrieved documents can be quite long and often contain a lot of redundant information, which can further exacerbate the problem of low inference efficiency.
@@
Inference efficiency is a key concern in practical applications of MLLMs. However, previous work, including RAVQA-v2, has focused on how to use as much knowledge as possible to improve the accuracy of answers while neglecting the fact that inference efficiency significantly declines as the number of input tokens increases.
@@
Therefore, we aim to design an innovative RAG framework based on MLLMs, which can utilize the information of retrieved knowledge in an effective and efficient manner to improve MLLMs’ inference efficiency for KB-VQA.
@@
Furthermore, RAVQA-v2 and many previous works (Luo et al., 2021; Lin & Byrne, 2022; Lin et al., 2022) on KB-VQA primarily focus on using textual documents as external knowledge sources, and there has been relatively less research on using multimodal documents as knowledge sources. However, multimodal documents are an important knowledge resource in real-world applications, which can provide knowledge in handling KB-VQA (Raffel et al., 2020; Hu et al., 2023b). What’s more, MLLMs inherently have the ability to directly comprehend multimodal knowledge. Therefore, we aim to explore the effects of multimodal documents in RAG applications based on MLLMs.
In this paper, we propose RACC, i.e. Retrieval-Augmented MLLMs with Compressed Contexts, an effective and efficient RAG framework for KB-VQA based on MLLMs. RACC consists of three phases, namely compression learning, information aggregation, and modulation generation. Our contributions can be summarized as follows:
  • RACC is the first work to integrate prompt compression technology with KB-VQA, proposing an innovative framework for efficient RAG on MLLMs. We identify four key issues in training RACC and propose four corresponding methods to address them.
@@
  • RACC achieves excellent performance comparable to many competitive baselines on two KB-VQA datasets at a very low cost, reaching a state-of-the-art (SOTA) performance of 63.92% on the OK-VQA dataset.
@@
  • RACC achieves outstanding enhancements in terms of time and space efficiency. For time efficiency, RACC can save 22.0-59.7% of inference latency compared to RAVQA-v2. For space efficiency, RACC supports pre-saving documents that occupy a large storage footprint in the form of compressed prompts to save disk space.
@@
  • RACC can be applied to various off-the-shelf MLLMs, but also can handle different knowledge sources such as textual documents and multimodal documents.
Figure 1: The structural framework of RACC. An image and a question are first input into the multimodal retriever to retrieve $K$ relevant documents. During the compression learning phase, the $K$ documents, image, and question are input into HyperMLLM to obtain their corresponding compressed prompts. In the information aggregation phase, the obtained compressed prompts are aggregated to form the document-based compressed prompts of vision and question. In the modulation generation phase, the output of the aggregation network is processed by a set of MLPs to obtain the KV cache for each layer of the downstream BaseMLLM. At the same time, the BaseMLLM receives the image and question as input and generates the final answer.
2 Related Work
2.1 Multimodal Large Language Models
@@
Multimodal large language models (MLLMs) bridge the gap between text and other modalities and unify understanding of different modalities. Although recent MLLMs (Alayrac et al., 2022; Li et al., 2023; Dai et al., 2023; Bai et al., 2023; Liu et al., 2023; Hu et al., 2024) have demonstrated excellent zero-shot performance in VQA tasks, they require the assistance of external knowledge sources for questions that require specialized domain knowledge.
In this work, we investigate two main architectures of MLLMs: encoder-decoder MLLMs including BLIP2-FlanT5XL (Li et al., 2023) and InstructBLIP-FlanT5XL (Dai et al., 2023), and the decoder-only MLLMs including InstructBLIP-Vicuna7B and miniCPM-v2 (Hu et al., 2024).
2.2 Knowledge-based Visual Question Answering
Knowledge-based Visual Question Answering (KB-VQA), anchored by the given images, works to answer questions that require external knowledge. KB-VQA is an important multimodal task and has garnered widespread attention.
@@ 2.2 Knowledge-based Visual Question Answering
Early KB-VQA works (Luo et al., 2021; Marino et al., 2021) train specialized VQA models with designated knowledge sources such as ConceptNet KnowledgeGraph (Speer et al., 2017), textual document sources like Google Search (Luo et al., 2021) or WikiData (Vrandečić & Krötzsch, 2014), and image sources like Google Image Search etc.. However, these methods often demonstrate limited performance.
@@
With the emergence of pretrained large language models (LLMs), they have become a focus of research in this field. A series of works utilize very large LLMs like GPT-3 (Brown et al., 2020), chatGPT to generate auxiliary knowledge or directly answer the questions (Gui et al., 2021; Lin et al., 2022; Yang et al., 2022; Shao et al., 2023; Hu et al., 2023a), achieving performance breakthroughs. Due to the high costs associated with using very large LLMs like GPT-3, and the fact that the knowledge within them can be outdated and incorrect, another line of work focuses on how to conduct retrieval-augmented generation (RAG) on smaller LLMs with various external knowledge bases for KB-VQA (Gui et al., 2021; Gao et al., 2022; Lin & Byrne, 2022; Lin et al., 2024a; Hu et al., 2023b). However, the above methods all face the same problem: they require converting images into textual descriptions such as captions, object tags, etc., so that LLMs can understand (Gui et al., 2021; Gao et al., 2022; Shao et al., 2023; Lin et al., 2024a), which may result in loss of critical visual information, but also significantly increases the number of input tokens, leading to a notable increase in inference latency.
@@
With the advent of multimodal large language models (MLLMs), the aforementioned problems have been perfectly solved. Recent works have made great progress by utilizing MLLMs. A line of works proposes to combine MLLMs and LLMs together (Khademi et al., 2023; Xenos et al., 2023; An et al., 2024; Liang et al., 2024, 2025). MM-Reasoner (Khademi et al., 2023) leverages vision APIs and rationales generated by GPT-4 to fine-tune MLLMs such as Flamingo. RAVQA-v2 (Lin et al., 2024a) is the first work to build a simple RAG framework on top of MLLMs. However, it suffers from low efficiency during the inference stage and doesn’t study the impact of multimodal documents on RAG applications of MLLMs.
2.3 Prompt Compression
@@
Given the inherent redundancy in natural language, prompt compression methods have been extensively studied to improve the efficiency of LLM inference. Prompt compression can be categorized into task-aware and task-agnostic methods. Since the generation of compressed prompts that perform well across diverse tasks is particularly challenging, we focus on the task-aware prompt compression paradigm.
An important line of work Jiang et al. (2023a); Pan et al. (2024); Jiang et al. (2023b) estimates the importance of the tokens within the original prompts by the information-based metrics and removes redundant tokens. Xu et al. (2024) trains a compressor model, which cuts redundant tokens in the passage based on the question.
@@
In addition to the above methods, which detect and remove inherent redundant tokens in long contexts at the natural language level, a series of works aims at compressing long contexts into parameters, leveraging the capability of LLMs to implicitly eliminate redundant information within long contexts. Mu et al. (2024) supposes that each prompt is composed of a task instruction part and a content part, and finetunes LLMs to compress the task instruction part into several gist tokens. Chevalier et al. (2023); Wang et al. (2024); Tack et al. (2024) proposed to compress long contexts into compact summary vectors, parameters of a Lora-module and KV Cache, respectively. Such kind of methods have already demonstrated significant research value and strong research potential. In this paper, we extend these methods to the KB-VQA task.
@@
RACC-hetero BLIP2-Vicuna7B GS 61.65
RACC-hetero InstructBLIP-Vicuna7B GS 63.92
@@
Table 1: Model Performance on the OK-VQA dataset. Knowledge source abbreviations: C: ConceptNet; CC: CC12M; V2: VQA-2; W: Wikipedia; WD: WikiData; WIT: Wikipedia Image-Text; GS: Google Search; GI: Google Images; ICE: In-context Examples; PNP: Plug-and-Play VQA captioner (Tiong et al., 2022). Please refer to Section 3.4 for the implications for RACC-homo and RACC-hetero. In the last two rows of results of the RACC-hetero, the hyperMLLM used is InstructBLIP-FlanT5XL.
3 Methods
In this section, we introduce our proposed RACC, i.e. Retrieval-Augmented MLLM with Compress Contexts. RACC can be divided into three phases: compression learning, information aggregation, and modulation generation, which will be detailed in the corresponding subsections. Based on a profound understanding of the task and an analysis of the shortcomings in RAVQA-v2, we identify four key issues that need to be addressed in training RACC and further propose four methods to address them, which is elaborated in Subsection 3.2 and 3.3.
3.1 Problem Setup
@@
A typical VQA dataset can be divided into three components: images, questions, and answers, which can be represented using the notation $\{v,q,a\}_{n}$. Following Lin et al. (2024a), we consider a realistic scenario of KB-VQA: an MLLM that takes an image $v_{i}$ and its related question $q_{i}$ as input, where the knowledge required to answer the question is supplied by external knowledge sources. In this paper, we study two important knowledge sources in real-world applications: multimodal documents and textual documents.
@@
We utilize an off-the-shelf frozen multimodal retriever to retrieve $K$ documents from the given knowledge source conditioned on the provided image and question. The $K$ retrieved documents is denoted as $\{d_{i}\}^{K}=\{d^{1}_{i},d^{2}_{i},\ldots,d^{K}_{i}\}$.
The confidence scores that the multimodal retriever outputs for each of the $K$ documents, i.e. the document retrieval scores, are denoted as $\{p_{i}\}^{K}=\{p^{1}_{i},p^{2}_{i},\ldots,p^{K}_{i}\}$. The MLLM needs to leverage these retrieved documents to provide the correct answer to the question $q_{i}$ based on $v_{i}$.
3.2 Phase 1: Compression Learning
@@
The first phase of RACC is to compress the retrieved document into soft prompts of a specified length. Phang et al. (2023) and Tack et al. (2024) introduce the idea of amortized-based meta-learning into the online learning task, which brings us inspiration. We utilize MLLMs of encoder-decoder architecture (such as BLIP2-FlanT5XL (Li et al., 2023), InstructBLIP-FlanT5XL (Dai et al., 2023), etc.) with a set of learnable prompts to compress the input information.
We denote the hyperMLLM by $\textsc{M}_{hyper}$ and represent the encoder and decoder within the hyperMLLM by $\textsc{Enc}_{hyper}$ and $\textsc{Dec}_{hyper}$, respectively. Each retrieved document is first fed into $\textsc{Enc}_{hyper}$. Subsequently, $\textsc{Dec}_{hyper}$ takes the output from the encoder and the learnable prompts $\theta_{d}$ as inputs, and the resulting output is the compressed prompts corresponding to that document. After such a compression process, the contextual information of each document is preserved in its compressed prompts. Consequently, in subsequent processing steps, the number of tokens that need to be processed for each document is reduced from the original token count of the document to the predefined length of $\theta_{d}$. The compressing process can be denoted as follows:
$\theta^{k}_{i}=\textsc{M}_{hyper}(d^{k}_{i},\theta_{d})=\textsc{Dec}_{hyper}(\textsc{Enc}_{hyper}(d^{k}_{i}),\theta_{d})$ (1)
$\theta^{k}_{i}$ denotes the compressed prompts of the document $d^{k}_{i}$.
We also compress the given image-question pair in a similar manner. Note that the predefined learnable prompts for compressing the image-question pairs are different from $\theta_{d}$, which we denote as $\theta_{vq}$, the compression process of the given image-question pair is denoted as follows:
$\begin{split}&\theta_{vq_{i}}=\textsc{M}_{hyper}(\textsc{concat}(v_{i},q_{i}),\theta_{vq})\end{split}$ (2)
We name this phase the compression learning phase, as the learnable prompts need to be trained to acquire the ability to compress document context information. We then identify two key issues in this phase and propose corresponding solutions, as described below:
Issue 1: How to initialize the learnable prompts?
Analysis: First, the lengths of the two sets of learnable prompts, i.e. $\theta_{d}$ and $\theta_{vq}$, are critical in the compression learning process. If the prompts are too short, the amount of compressed semantic information that they can retain will be insufficient. In contrast, if the prompts are too long, the difficulty of training increases significantly.
We suggest determining the lengths of two sets of learnable prompts by experiments. To determine the optimal lengths of these prompts, we conducted a series of comparative experiments. We treat the lengths of $\theta_{d}$ and $\theta_{vq}$ as hyperparameters and systematically explore their impact on the compression learning process.
Additionally, the initialization weights of the learnable prompts also play a crucial role in the hyperMLLM’s ability to compress inputs, especially in the early training stage.